. Find the square of the following numbers.  vi. 46 
,jg,kymtdjrdtxhyrfjkhyktjuryhrdjyk,y

Grade:12th pass

1 Answers

Harshit Singh
2 years ago
Dear Student


vi. (46)2
=(50-4)2
= (50)2+ (4)2- 2×50×4 =2500+16–400
= 2116


Thanks